Status Report

ISS MER Management’s Daily Notes Monday, June 13, 2001 – 12:00 pm

By SpaceRef Editor
June 13, 2001
Filed under , ,

The following notes highlight the ISS activities within the last 24 hours that were discussed during the morning NASA and Boeing management teleconferences, as well as the daily MER Management & Subsystem Tag-Up. These notes are not intended to provide the current configuration of the ISS systems, nor are they intended to capture all of the events that have occurred in the past 24-hour period or that are planned for the next 24-48-hour period. For a detailed overview of the last 24-hour period and upcoming ISS events, refer to the “0800 Report” and “Daily Report”, published by MOD on its Mission Control Center (MCC) website at (select “Realtime Mission,” then “ISS Msgs”).

  • Attitude: The ISS successfully transitioned from XPOP to XVV around 6:45am this morning. The TEA is ~ 0,-10,0.

  • Channel 4B BGA Testing: The 4B Solar Array BGA was rotated continuously for several orbits yesterday to obtain additional data to support investigation of the high-current anomaly. The test called for certain actions to be taken whenever the current exceeded 0.5-amp for greater than 10-min. However, during that time, the current remained around 0.3-amp and never exhibited the high-current anomaly. The structures and mechanisms team, as well as the ART, is currently evaluating the data to determine whether the 4B array should be placed in rate mode for an extended period of time now that the ISS is back in XVV. The 2B array is currently in rate mode.

  • ESP-1 Heater Issue: During the ART conducted yesterday, June 12, the thermal team presented its thermal analysis data for the DCSU and PFSC. The ART was informed that the DCSU reached -57-deg F, which exceeds the ORU acceptance thermal limit of -45 deg F remained above the -65 deg F qualification limit. The PFCS reached a temperature of -74 deg F, which exceeded both the -45 deg F acceptance and -65 deg F qualification limit. Analysis is still ongoing to determine what these temperature environments could have done to the functionality/reliability of the ORUs, as well as what capabilities we have (e.g., spare ORUs, flight support equipment, manifest room on 7A.1 or beyond) to remove and replace the spare ORUs on-orbit. As part of the investigation, it was also determined that there may be a problem with one of the DCSU heaters on ESP-1. There are 4 heaters, but only 3 came on when the RPCs were closed. This could indicate a failed heater. Investigation is ongoing, but preliminary indications are that the 3 heaters that are functioning will provide sufficient margin for maintaining ESP ORUs within limits. An updated status will be presented to the VSIP on Thursday, June 14.

  • SSRMS: Software Patch #3, which will be used to perform additional diagnostic testing of the 1553 bus Remote Terminal (RT) addressing anomaly, will be uploaded to the ISS this afternoon (tentatively around 12:00-12:30pm). Shortly thereafter, additional diagnostic testing will be performed to further isolate the problem. The testing will involve power cycling the redundant string in an attempt to re-introduce the transmit/receive error. Once the testing is complete, the patch will be removed so that the SSSRM airlock dry run can be performed tomorrow, which will be done using the SSRMS redundant string (as planned for Flight 7A).

  • XPOP Plasma Characterization Testing: Testing to characterize the plasma environment and vehicle charging potential while in XPOP was completed yesterday, June 12. The emission current from PCU 2, which is in discharge mode, was measured to gauge the vehicle potential while array shunting was performed at various times throughout several orbits. The PCU Tiger Team is currently evaluating the data as part of its efforts to determine specific shunting and/or array pointing constraints, if any, that are required to support EVAs performed when the ISS is in XPOP.

  • EEPROM Bit Flip: A health flag status was annunciated yesterday for the Node 1 N1-4BA RPCM. The block in which the health flag was had a flipped data bit that was previously refreshed on GMT day 102.

  • MCA: Yesterday, it was reported that the MCA went to idle mode when it experienced a Purge Time Override Failure, an anomaly that can occur when the MCA is transitioning from Rapid Sampling to Auto Sequence mode. (Note that it has not been confirmed that this is what caused the Purge Time Override Failure to occur, so a new IFI will be generated when access to the server is restored). Since vacuum was lost for an extended period of time, the MCA will need to be pumped out to bring it back online. Because of the crew time required to support this activity, tentative plans are to hold-off on pumping out the MCA and bringing it back online until after the MCA components are removed and replaced, which is scheduled for approximately 7-day prior to 7A.

  • NCS PCU Patch: Upload of a NCS PCU Patch to improve synchronization between the NCS and PCU to prevent inadvertent PCU shutdown is scheduled for today.

  • O2 Repress: ISS repress was nominally performed using O2 from the Progress.

  • Ku-Band Auto-track Anomaly: Prior to performing a TDRS handover on the Ku-Band system last night, the Ku-Band was operating nominally. The command to configure Ku-Band back to auto-track mode was given and accepted. As expected, the system transitioned to open loop slew and began the spiral acquisition sequence. The TD171 forward link signal was found during the spiral search and the system transitioned to acquisition lock state with a PWRL of -47 dB. The PLC state transitioned from reset to normal momentarily, but then transitioned back to reset, so the Ku-Band transmitter output power remained at zero. The Ku-Band Antenna was neither near a mask nor did the PWRL value drop below -51 dB at any time after the initial acquisition sequence. The system remained in auto-track mode and continued to track TD171 forward link until the end of the TDRS event. The PLC state remained in reset for the entire event. During the next Ku-Band opportunity, system was commanded back to auto-track and it began to function properly. A new IFI is being established to investigate this issue.

  • BP/ECG Power Trip: After plugging in the Blood Pressure/Electro-Cardiogram machine into a UOP this morning, the UOP tripped. Initial troubleshooting involved replacing the power cable to the BP/ECG to determine whether the power trip occurred because of the BP/ECG or was attributable to the UOP. When the BP/ECG was tested with the new cable, the UOP tripped again. Subsequently, the BP/ECG was tried in another UOP, and that UOP tripped when activation of the BP/ECG was attempted, indicating that the problem most likely lies with the BP/ECG and not the UOP. A new IFI is being established to investigate this issue.

  • Progress Thruster Testing: ISS attitude control was successfully handed over the Russian Segment to support Progress Thruster Firing planned by the Russians. Following completion of the thruster firing tests, attitude control will be handed back over to the U.S. Segment. From a Russian perspective, the Progress would then be used to control ISS pitch and yaw, with the SM retaining roll control.

    Today’s Events (GMT 2001/164):

  • Upload diagnostic software patch to the SSRMS
  • Russian testing of the Progress Vehicle thrusters

    Upcoming Events for GMT 2001/165-166 (i.e., Thursday and Friday):

  • SM Battery R&R (swapping SM batteries in Locations 2 and 6) – Thursday, June 14
  • SSRMS Operations (i.e., Airlock Dry Run installation procedure) on Redundant String, Thursday, June 14

    Other Upcoming Events:

  • SSRMS ART — June 13, 11am, MER Conference Room
  • ESP-1 Heater ORU Presentation to VSIP – June 14, Building 4S, Room 3419, ~3pm
  • Thursday, June 14 – MCA ART, 1pm (tentative – may be canceled)
  • Thursday, June 14 – Smoke Detector ART, location and time TBD
  • Friday, June 15 – Vacuum System ART, MER Conference Room, 2:00 pm
  • Tuesday, June 19 – Upgrade to new web-based IFI Database, ~3pm

  • SpaceRef staff editor.